# Talha Bozkus and Urbashi Mitra

### v2

@@ V-B Deterministic Analysis
This proposition provides insight into the behavior of the joint $Q$-function updates over time. As more iterations are performed, the updates to the joint $Q$-function become increasingly stable, indicating the stabilization of the overall learning process. This result applies to joint $Q$-functions and thus provides generalization over to the single-agent case [6], which is restricted to individual $Q$-functions. We leverage the stability of each MEMQ agent from [6] to derive the stability of the joint system. The proof is given in Appendix A.
@@
This result is valid for any $t$; thus, it can be used as a stopping condition for Algorithm 1. Moreover, by appropriately choosing the parameters $u$ and $K_{i}$, we can effectively control the speed of convergence and the accuracy of the $Q$-function updates. This result reflects the sample complexity of Algorithm 1 to achieve $\beta$-convergence – this result is different than the standard sample complexity, which considers $\bar{Q}_{t}-\bar{Q}^{*}$ yet we herein consider $\bar{Q}_{t}-\bar{Q}_{t-1}$. This result considers the aggregated performance of all agents and requires more iterations to match the sample complexity of the single-agent case [5].
@@
Let $\bar{s}$ be an uncoordinated joint state. If $|\epsilon^{(n)}_{i,t-1}(s_{i},a_{i})|\leq|\epsilon^{(n)}_{i,t}(s_{i},a_{i})|$ for all $n,t,i$, then:
@@ V-C Analysis of State Misdetection Probability
where $I_{i,t}$ and $I_{i,t,\text{true}}$ are the noisy and true ARSS measurement by TXi due to the other transmitters at time $t$ and $n_{t}$ is additive zero-mean Gaussian noise as: $n_{t}\sim N(0,\sigma_{u}^{2})$ in uncoordinated states and $n_{t}\sim N(0,\sigma_{c}^{2})$ in coordinated states. In uncoordinated states, transmitters are often far apart from each other, leading to higher measurement uncertainty due to a lack of information exchange. Conversely, in coordinated states, closer proximity and synchronization reduce this uncertainty. Hence, we assume $\sigma_{u}\gg\sigma_{c}$.
@@
$P_{\text{mis}}\in\Bigg{[}Q\left(\frac{\Delta_{I}}{\sigma_{c}}\right)$ $\frac{|\mathcal{S}_{C}|}{|\mathcal{S}|}+Q\left(\frac{-\Delta_{I}}{\sigma_{u}}\right)\frac{|\mathcal{S}_{U}|}{|\mathcal{S}|},$
$Q\left(\frac{-\Delta_{I}}{\sigma_{c}}\right)\frac{|\mathcal{S}_{C}|}{|\mathcal{S}|}+Q\left(\frac{\Delta_{I}}{\sigma_{u}}\right)\frac{|\mathcal{S}_{U}|}{|\mathcal{S}|}\Bigg{]},$ (18)
where $Q(z)=\frac{1}{2}\operatorname{erfc}\left(\frac{z}{\sqrt{2}}\right)$, and $\operatorname{erfc}$ is the complementary error function, and
@@
The lower bound represents the fundamental limit on detection accuracy due to noise in ARSS measurements, reflecting an unavoidable misdetection error. The upper bound indicates the maximum misdetection probability, representing the worst-case scenario where the system is most vulnerable to errors from noisy ARSS measurements. We note that the lower and upper bounds are complementary and sum to 1. Thus, an increase in the lower bound directly results in a decrease in upper bound, and vice versa. For example, as $\sigma_{c}$ increases, the lower bound increases, indicating a higher minimum error probability. On the other hand, the upper bound decreases because higher noise in coordinated states causes the ARSS measurements to resemble those of uncoordinated states. As a result, the likelihood of misdetecting all states in the worst-case scenario decreases. When $|\mathcal{S}_{C}|\to|\mathcal{S}_{U}|$ and $\sigma_{C}\to\sigma_{U}$, both bounds converge to 0.5, indicating that the system can no longer distinguish between coordinated and uncoordinated states, leading to random detection. The proof is given in Appendix B.
@@
The bounds on the probability of misdetection can be generalized to any number of agents $N_{T}>2$ using Taylor series approximation as follows: (See [46] for proof)
$P_{mis}\in\Bigg{[}$ $\Big{(}1-\sum_{i=1}^{N_{T}}Q\Big{(}\frac{I^{*}_{thr}-I_{i,t,true}}{\sigma_{c}}\Big{)}\Big{)}\frac{|\mathcal{S}_{c}|}{|\mathcal{S}|}+$
$\hskip 40.0ptQ\Big{(}\frac{I^{*}_{thr}-I_{i^{\prime},t,true}}{\sigma_{u}}\Big{)}\frac{|\mathcal{S}_{u}|}{|\mathcal{S}|},$
$\sum_{i=1}^{N_{T}}Q\Big{(}\frac{I^{**}_{thr}-I_{i,t,true}}{\sigma_{u}}\Big{)}\frac{|\mathcal{S}_{u}|}{|\mathcal{S}|}+$
$\hskip 40.0pt\Big{(}1-Q\Big{(}\frac{I^{**}_{thr}-I_{i^{\prime},t,true}}{\sigma_{c}}\Big{)}\Big{)}\frac{|\mathcal{S}_{c}|}{|\mathcal{S}|}\Bigg{]},$ (20)
where $I^{*}_{thr}$ and $I^{**}_{thr}$ are two distinct constants as a (non-linear) function of $N_{T}$, $|\mathcal{S}_{c}|$, $|\mathcal{S}_{u}|$, $\sigma_{c}$, $\sigma_{u}$ and the minimum true ARSS measurement among all agents $I_{i^{\prime},t,true}=\underset{i=1,\dots,N_{T}}{\min}\,I_{i,t,true}$.
This result is derived using a second-order Taylor series approximation of the Gaussian $Q(x)$ function around $x=0$ and hence is numerically accurate when the threshold $I_{thr}$ is close to the true ARSS measurements $I_{i,t,true}$ or when noise variances $\sigma_{c}$ and $\sigma_{u}$ are large. It is possible to employ a higher-order Taylor approximation for a more precise bound, yet it becomes harder to derive a closed-form expression. The lower and upper bounds herein are complementary (i.e., one increases while the other one decreases) as in the two-agent case. We will numerically analyze the effect of individual parameters on the tightness of the bound.
@@ VI Numerical Results
In this section, we consider a variety of performance metrics to assess the accuracy and complexity performance of Algorithm 1. The following parameters are constant across all simulations: $T=10^{5}$, $l=30$, $\gamma=0.95$, $\alpha_{t}=\frac{1}{1+\nicefrac{{t}}{{1000}}}$, $\epsilon_{t}=\max(0.99^{t},0.01)$, $u_{t}=1-e^{\nicefrac{{-t}}{{1000}}}$ (if constant $u=0.5$ ), $L=100$, $\Delta_{L}=2$, $r_{j}\sim\operatorname{unif}\left[\frac{L}{2},\frac{3L}{4}\right]$, $\Delta_{I}=50$, $\sigma=1$, $P_{i}=10^{-2}$, $\beta_{1}=\beta_{2}=\beta_{3}=\frac{1}{3}$. The parameters $I_{max}$ and $I_{min}$ are set to the max and min ARSS measurements for each simulation. The parameters $\lambda_{i,n}$ are estimated numerically as in [5]. $I_{thr}$ is chosen to be the minimizer of $P_{mis}$ from Proposition 4 and Corollary 4. $|\mathcal{S}_{c}|$ and $|\mathcal{S}_{u}|$ are determined using $I_{thr}$. $K_{i}$ is selected proportional to the state-space size as in [5], and remain constant for all $i$. The values of $\sigma_{c}$ and $\sigma_{u}$ are chosen such that $\sigma_{u}\gg\sigma_{c}$ and the assumption in Proposition 1 is satisfied. Parameters of joint state-estimation $\bm{\mu}$, $\bm{\Sigma}$, $\delta_{0}$ and $\mathbf{b}_{i,0}$ are determined as in [44]. Simulations are averaged over 100 independent simulations.
@@ VI-A Policy and $Q$-function accuracy
where $\bar{Q}^{*}$ is the optimal joint $Q$-functions, and $\bar{Q}$ is the joint output $Q$-functions of Algorithm 1 at time $t$. We evaluate our algorithm against several fully cooperative tabular and deep MARL algorithms as shown in Table III. These algorithms can be used with our wireless network setting and are chosen such that different MARL algorithms with different characteristics are represented.
@@
APE results across an increasing number of transmitters and base stations $(N_{T},N_{B})$ are shown in Fig.3a. Our algorithm offers 35% less APE than CTDE, 45% less APE than DNA, and 55% less APE than fully decentralized algorithms, particularly across large $(N_{T},N_{B})$. These APE gains can be attributed to the following reasons: (i) exchanging local ARSS information through the leader can effectively capture coordination among agents. (ii) Leveraging synthetic environments through MEMQ helps agents improve their exploration capabilities [5]. (iii) MEMQ exploits structural similarities among multiple environments, providing an advantage over neural network-based algorithms that do not utilize such properties. (iv) Optimizing $I_{thr}$, by minimizing the state misdetection probability, achieves the optimal $|\mathcal{S}_{c}|/|\mathcal{S}_{u}|$ ratio, which combines coordination benefits in coordinated states and complexity gains in uncoordinated states. The centralized algorithm achieves slightly better APE than our algorithm, as expected since it assumes full observability among agents. However, the performance difference is minimal (only less than 10%) because (i) communication and information exchange among agents only in coordinated states are sufficient to capture coordination, and (ii) periodically resetting the belief state prevents error accumulation and yields accurate joint state estimates. We also note that APE improvements increase as the wireless network grows, demonstrating the scalability of the proposed algorithm.
@@
Fig.3b shows the AQD results over iterations with $(N_{T},N_{B})=(5,3)$. $Q$-functions of all algorithms are initialized to very small random but identical values (thus AQD of all algorithms start at the same value). AQD values for all algorithms decrease over time with varying decay rates, convergence behavior, and convergence points. This result demonstrates that our algorithm converges numerically to the optimal $Q$-functions, with a convergence rate 25% faster than CTDE and DNA algorithms, and 35% faster fully decentralized algorithms. This result follows from the fast exploration and convergence properties of the single-agent MEMQ and the purely decentralized nature of agents in uncoordinated states. This result also shows the accuracy and convergence of the Bayesian joint state estimation technique. Our algorithm has a similar converge speed to the centralized algorithm, which is consistent with Fig.3a. We also observe that decentralized algorithms (e.g., HQ and I2Q) and deep RL algorithms (e.g., DQN) may not always converge to the optimal $Q$-functions. We underline that algorithms have similar trends in Fig.3a and Fig.3b (e.g. HQ has the largest AQD as it does not converge, which also leads to a very large APE in the limit).
@@
$r_{j}$ $\operatorname{unif}[x,3x],x\in\{\frac{L}{5}$,$\frac{L}{4}$,$\frac{L}{3}$,$\frac{L}{2}$} 8% 8%
@@ VI-B Complexity
In order to evaluate the complexity of Algorithm 1, we first consider computational runtime complexity – the time required for convergence with optimized parameters on the same hardware. Though runtime varies by hardware, we focus on the relative complexity between different algorithms. Some algorithms may not converge to the optimal $Q$-functions; in such cases, we terminate the algorithm once it converges to a stable point. Fig.4 shows the runtime results across increasing $(N_{T},N_{B})$. While all algorithms show an increase in runtime as the wireless network grows, our algorithm has a relatively small increase. In particular, our algorithm can achieve up to 50% less runtime than fully centralized, CTDE and DNA algorithms, particularly across large $(N_{T},N_{B})$. This efficiency stems from (i) a fully decentralized approach in uncoordinated states, which eliminates unnecessary coordination and information exchange among agents; (ii) minimal information sharing in coordinated states through only the leader transmitter, which avoids expensive and redundant communication between all agents, and (iii) leveraging fast exploration capabilities of digital cousins and fast convergence of single-agent MEMQ. While decentralized algorithms assume full independence between agents without any coordination or communication, and hence have the smallest runtime, our algorithm has only %10 larger runtime than the decentralized algorithms. We underline that this result is consistent with Fig.3b as plotting the AQD values of algorithms over iterations for different $(N_{T},N_{B})$ values yields similar plots to Fig.4 (i.e. similar algorithm orders).
One can also consider the algorithm sample complexity – the number of training samples (or interactions with the environment) needed to achieve good policies. To this end, we can determine the time index at which the AQD in (22) falls below some threshold. This index is proportional to the sample complexity, as the total number of training samples obtained by iteration $t$ is $t$ x $l$. Fig. 3b provides an example where our algorithm achieves an AQD of less than 1 in 22000 iterations, while other algorithms need at least 40000 iterations, resulting in a 45% reduction in sample complexity. This improvement becomes greater for smaller thresholds (e.g. 55% reduction in sample complexity to reach AQD < 0.5).
@@
Extensive simulations also show that both the computational runtime and sample complexity of our algorithm are primarily dominated by the slowest agent (i.e. the slowest MEMQ). This is expected, as the joint system cannot converge until all individual systems do. Therefore, optimizing the slowest agent is an effective strategy for enhancing the overall complexity of the joint system.
The memory complexity of Algorithm 1 (i.e., storing local and global $Q$-functions for each agent), is not explicitly addressed in this paper; however, it can be handled by employing the structural state-aggregation algorithm presented in [9], which is successfully employed for wireless networks in [6].

### v3

@@ V-B Deterministic Analysis
This proposition provides insight into the behavior of the joint $Q$-function updates over time in uncoordinated states. As more iterations are performed, the updates to the joint $Q$-function become increasingly stable, indicating the stabilization of the overall learning process. This result applies to joint $Q$-functions and thus provides generalization over to the single-agent case [5], which is restricted to individual $Q$-functions. We leverage the stability of each agent from [5] to establish joint stability, since in uncoordinated states agents act independently and the single-agent MEMQ results apply directly to each agent.
@@
This result is valid for any $t$; thus, it can be used as a stopping condition for Algorithm 1. Moreover, by appropriately choosing the parameters $u$ and $K_{i}$, we can effectively control the speed of convergence and the accuracy of the $Q$-function updates. This result reflects the sample complexity of Algorithm 1 to achieve $\beta$-convergence, which is different than the standard sample complexity, which considers $\bar{Q}_{t}-\bar{Q}^{*}$ yet we herein consider $\bar{Q}_{t}-\bar{Q}_{t-1}$. This result considers aggregated performance of all agents and requires more iterations to match the sample complexity of the single-agent case [4].
@@
Let $\bar{s}$ be an uncoordinated joint state. If $|\epsilon^{(n)}_{i,t-1}(s_{i},a_{i})|\geq|\epsilon^{(n)}_{i,t}(s_{i},a_{i})|$ for all $n,t,i$, then:
@@ V-C Analysis of State Misdetection Probability
where $I_{i,t}$ and $I_{i,t,\text{true}}$ are the noisy and true ARSS measurement by TXi due to the other transmitters at time $t$ and $n_{t}$ is additive zero-mean Gaussian noise as: $n_{t}\sim N(0,\sigma_{u}^{2})$ in uncoordinated states and $n_{t}\sim N(0,\sigma_{c}^{2})$ in coordinated states. In uncoordinated states, transmitters are often far apart from each other, leading to higher measurement uncertainty due to a lack of information exchange. Conversely, in coordinated states, closer proximity and synchronization reduce this uncertainty. Hence, we assume $\sigma_{u}\gg\sigma_{c}$. Additionally, we define the probability of state misdetection as:
$P_{\mathrm{mis}}$ $=P(\text{detect U}\mid\text{actual C})P(\text{actual C})$
$\quad+P(\text{detect C}\mid\text{actual U})P(\text{actual U}),$ (21)
[…]
Using the noisy ARSS model (17), we can derive a lower bound on $P_{\mathrm{mis}}$ as a function of $I_{\mathrm{thr}}$ (see the supplementary material), which represents the unavoidable state misdetection error. Hence, varying $I_{\mathrm{thr}}$ directly impacts the accuracy of state detection. If $I_{\mathrm{thr}}$ is set too high, most states will be misclassified as uncoordinated, even when coordination is beneficial. This leads to high policy error due to missed opportunities for joint optimization, but communication and joint updates are rarely triggered, resulting in minimal runtime and overhead. Conversely, setting $I_{\mathrm{thr}}$ too small causes nearly all states to be classified as coordinated, yielding excessive communication with the leader and frequent joint $Q$-function updates even when unnecessary. This may improve policy accuracy, yet incurs higher computational overhead. To this end, we choose $I_{\mathrm{thr}}$ to minimize the lower bound on $P_{\mathrm{mis}}$. The optimal $I_{\mathrm{thr}}$ admits a closed-form expression as given in the supplementary material, and yields the following bounds on $P_{\mathrm{mis}}$.
@@
$P_{\text{mis}}\in\Bigg{[}\mathcal{Q}\left(\frac{\Delta_{I}}{\sigma_{c}}\right)$ $\frac{|\mathcal{S}_{C}|}{|\mathcal{S}|}+\mathcal{Q}\left(\frac{-\Delta_{I}}{\sigma_{u}}\right)\frac{|\mathcal{S}_{U}|}{|\mathcal{S}|},$
$\mathcal{Q}\left(\frac{-\Delta_{I}}{\sigma_{c}}\right)\frac{|\mathcal{S}_{C}|}{|\mathcal{S}|}+\mathcal{Q}\left(\frac{\Delta_{I}}{\sigma_{u}}\right)\frac{|\mathcal{S}_{U}|}{|\mathcal{S}|}\Bigg{]},$ (22)
where $\mathcal{Q}(z)=\frac{1}{2}\operatorname{erfc}\left(\frac{z}{\sqrt{2}}\right)$ and
@@
The lower bound represents the fundamental limit on detection accuracy due to noise in ARSS measurements, reflecting an unavoidable misdetection error. The upper bound indicates the maximum misdetection probability, representing the worst-case scenario where the system is most vulnerable to errors from noisy ARSS measurements. We note that the lower and upper bounds are complementary and sum to 1. Thus, an increase in the lower bound directly results in a decrease in upper bound, and vice versa. For example, as $\sigma_{c}$ increases, the lower bound increases, indicating a higher minimum error probability. On the other hand, the upper bound decreases because higher noise in coordinated states causes the ARSS measurements to resemble those of uncoordinated states. As a result, the likelihood of misdetecting all states in the worst-case scenario decreases. When $|\mathcal{S}_{C}|\to|\mathcal{S}_{U}|$ and $\sigma_{C}\to\sigma_{U}$, both bounds converge to 0.5, indicating that the system can no longer distinguish between coordinated and uncoordinated states, leading to random detection.
@@
The bounds on the probability of misdetection can be generalized to any number of agents $N_{T}>2$ using Taylor series approximation as follows:
$P_{mis}\in\Bigg{[}$ $\Big{(}1-\sum_{i=1}^{N_{T}}\mathcal{Q}\Big{(}\frac{I^{l}_{thr}-I_{i,t,true}}{\sigma_{c}}\Big{)}\Big{)}\frac{|\mathcal{S}_{c}|}{|\mathcal{S}|}+$
$\hskip 40.0pt\mathcal{Q}\Big{(}\frac{I^{l}_{thr}-I_{i^{\prime},t,true}}{\sigma_{u}}\Big{)}\frac{|\mathcal{S}_{u}|}{|\mathcal{S}|},$
$\sum_{i=1}^{N_{T}}\mathcal{Q}\Big{(}\frac{I^{u}_{thr}-I_{i,t,true}}{\sigma_{u}}\Big{)}\frac{|\mathcal{S}_{u}|}{|\mathcal{S}|}+$
$\hskip 40.0pt\Big{(}1-\mathcal{Q}\Big{(}\frac{I^{u}_{thr}-I_{i^{\prime},t,true}}{\sigma_{c}}\Big{)}\Big{)}\frac{|\mathcal{S}_{c}|}{|\mathcal{S}|}\Bigg{]},$ (24)
where $I^{l}_{thr}$ and $I^{u}_{thr}$ are two distinct constants as a (non-linear) function of $N_{T}$, $|\mathcal{S}_{c}|$, $|\mathcal{S}_{u}|$, $\sigma_{c}$, $\sigma_{u}$ and the minimum true ARSS measurement among all agents $I_{i^{\prime},t,true}=\underset{i=1,\dots,N_{T}}{\min}\,I_{i,t,true}$.
This result is derived using a second-order Taylor series approximation of the Gaussian $\mathcal{Q}(x)$ function around $x=0$ and hence is numerically accurate when the threshold $I_{thr}$ is close to the true ARSS measurements $I_{i,t,true}$ or when noise variances $\sigma_{c}$ and $\sigma_{u}$ are large. It is possible to employ a higher-order Taylor approximation for a more precise bound, yet it becomes harder to derive a closed-form expression. The lower and upper bounds herein are complementary (i.e., one increases while the other one decreases) as in the two-agent case. We will numerically analyze the effect of individual parameters on the tightness of the bound.
@@ VI Numerical Results
In this section, we consider a variety of performance metrics to assess the accuracy and complexity performance of Algorithm 1. The following parameters are used: $T=10^{5}$, $l=30$, $\gamma=0.95$, $\alpha_{t}=\frac{1}{1+\nicefrac{{t}}{{1000}}}$, $\zeta_{t}=\max(0.99^{t},0.01)$, $u_{t}=1-e^{\nicefrac{{-t}}{{1000}}}$ (or $u=0.5$ ), $L=100$, $\Delta_{L}=2$, $r_{j}\sim\operatorname{unif}\left[\frac{L}{2},\frac{3L}{4}\right]$, $\Delta_{I}=50$, $\sigma=1$, $P_{i}=10^{-2}$, $\beta_{1}=\beta_{2}=\beta_{3}=\frac{1}{3}$. The parameters $I_{max}$ and $I_{min}$ are set to the max and min ARSS measurements for each simulation. The parameters $\lambda_{i,n}$ are estimated numerically as in [4]. $I_{thr}$ is chosen to be the minimizer of the lower bound on $P_{mis}$ from Proposition 4 and Corollary 4. $K_{i}$ is selected proportional to the state-space size as in [4], and remains constant for all $i$. The values of $\sigma_{c}$ and $\sigma_{u}$ are chosen such that $\sigma_{u}\gg\sigma_{c}$ and the assumption in Proposition 4 is satisfied. Simulations are averaged over 100 independent simulations.
@@ VI-A Policy and $Q$-function accuracy
where $\bar{Q}^{*}$ is the optimal joint $Q$-functions, and $\bar{Q}$ is the joint output $Q$-functions of Algorithm 1 at time $t$. We evaluate our algorithm against several fully cooperative tabular and deep MARL algorithms as shown in Table III. $\bar{Q}^{\star}$ and $\bar{\pi}^{\star}$ are computed via centralized $Q$-learning by treating all agents as a single composite agent with full joint information. This is used only for benchmarking and not required in practice.
@@
APE results across an increasing number of transmitters and base stations $(N_{T},N_{B})$ are shown in Fig.3a. Our algorithm offers 40% less APE than CTDE, 45% less APE than DNA, and 60% less APE than fully decentralized algorithms, particularly across large $(N_{T},N_{B})$. These APE gains can be attributed to the following reasons: (i) exchanging local ARSS information through the leader can effectively capture coordination among agents. (ii) Leveraging synthetic environments through MEMQ helps agents improve their exploration capabilities [4]. (iii) MEMQ exploits structural similarities among multiple environments, providing an advantage over neural network-based algorithms that do not utilize such properties. (iv) Optimizing $I_{thr}$, by minimizing the state misdetection probability, achieves the optimal $|\mathcal{S}_{c}|/|\mathcal{S}_{u}|$ ratio, which combines coordination benefits in coordinated states and complexity gains in uncoordinated states. The centralized algorithm achieves slightly better APE than our algorithm, as expected since it assumes full observability among agents. However, the performance difference is minimal (only less than 10%) because (i) communication and information exchange among agents only in coordinated states are sufficient to capture coordination, and (ii) periodically resetting the belief state prevents error accumulation and yields accurate joint state estimates. We also note that APE improvements increase as the wireless network grows, demonstrating the scalability of the proposed algorithm.
Our algorithm also supports highly heterogeneous users. We illustrate this with a 6-user system where each user’s parameters are randomly drawn from these sets: mobility levels $\{1\Delta_{L},2\Delta_{L},3\Delta_{L}\}$ steps per time slot; number of antennas $\{1,2,3,4\}$; transmission powers $\{0.1\,\text{W},0.25\,\text{W},0.5\,\text{W}\}$; and different coverage areas with radius $\{50\Delta_{L},75\Delta_{L},100\Delta_{L}\}$. Numerical results show that our algorithm reduces APE by 25% compared to CTDE, 40% compared to DNA, 45% compared to fully decentralized algorithms.
@@
Fig.3b shows the AQD results over iterations with $(N_{T},N_{B})=(5,3)$. $Q$-functions of all algorithms are initialized to very small random but identical values (thus AQD of all algorithms start at the same value). AQD values for all algorithms decrease over time with varying decay rates, convergence behavior, and convergence points. This result demonstrates that our algorithm converges numerically to the optimal $Q$-functions, with a convergence rate 30% faster than CTDE and DNA algorithms, and 40% faster fully decentralized algorithms. This result follows from the fast exploration and convergence properties of the single-agent MEMQ and the purely decentralized nature of agents in uncoordinated states. This result also shows the accuracy and convergence of the Bayesian joint state estimation technique. Our algorithm has a similar converge speed to the centralized algorithm, which is consistent with Fig.3a. We also observe that decentralized algorithms (e.g., HQ and I2Q) and deep RL algorithms (e.g., DQN) may not always converge to the optimal $Q$-functions. We underline that algorithms have similar trends in Fig.3a and Fig.3b (e.g. HQ has the largest AQD as it does not converge, which also leads to a very large APE in the limit).
@@
$r_{j}$ $\operatorname{unif}[x,3x],x\in\{\frac{L}{5}$$\frac{L}{4}$$\frac{L}{3}$$\frac{L}{2}$} 8% 8%
@@ VI-B Complexity
In order to evaluate the complexity of Algorithm 1, we first consider computational runtime complexity – the time required for convergence with optimized parameters on the same hardware. Though runtime varies by hardware, we focus on the relative complexity between different algorithms. Some algorithms may not converge to the optimal $Q$-functions; in such cases, we terminate the algorithm once it converges to a stable point. Fig.4 shows the runtime results across increasing $(N_{T},N_{B})$. While all algorithms show an increase in runtime as the wireless network grows, our algorithm has a relatively small increase. In particular, our algorithm can achieve up to 45% less runtime than fully centralized, CTDE and DNA algorithms, particularly across large $(N_{T},N_{B})$. This efficiency stems from (i) a fully decentralized approach in uncoordinated states, which eliminates unnecessary coordination and information exchange among agents; (ii) minimal information sharing in coordinated states through only the leader transmitter, which avoids expensive and redundant communication between all agents, and (iii) leveraging fast exploration capabilities of digital cousins and fast convergence of single-agent MEMQ. While decentralized algorithms assume full independence between agents without any coordination or communication, and hence have the smallest runtime, our algorithm has only 10% larger runtime than the decentralized algorithms. We underline that this result is consistent with Fig.3b as plotting the AQD values of algorithms over iterations for different $(N_{T},N_{B})$ values yields similar plots to Fig.4 (i.e. similar algorithm orders).
One can also consider the algorithm sample complexity – the number of training samples (or interactions with the environment) needed to achieve good policies. To this end, we can determine the time index at which the AQD in (26) falls below some threshold. This index is proportional to the sample complexity, as the total number of training samples obtained by iteration $t$ is $t$ x $l$. Fig. 3b provides an example where our algorithm achieves an AQD of less than 1 in 24000 iterations, while other algorithms need at least 40000 iterations, resulting in a 40% reduction in sample complexity. This improvement becomes greater for smaller thresholds (e.g. 50% reduction in sample complexity to reach AQD < 0.5).
@@
Extensive simulations also show that both the computational complexity of our algorithm are primarily dominated by the slowest agent (i.e. the slowest MEMQ). This is expected, as the joint system cannot converge until all individual systems do. Thus, optimizing the slowest agent is an effective strategy for enhancing the overall complexity of the joint system. The memory complexity of Algorithm 1 (storing local and global $Q$-functions for each agent), is not explicitly addressed in this paper; however, it can be handled by employing the structural state-aggregation algorithm which is presented in [6] and successfully employed for wireless networks in [5].